OpenStudy (anonymous):

Evaluate the expression by using the given value of the variables. y ÷ 2 • 4 – 10 ÷ x + 6 where x = 5 and y = 16 A. 3 B. 36 C. 21 D. –3

OpenStudy (alexandervonhumboldt2):

8*4-2+6=32-2+6=36

whitemonsterbunny17 (whitemonsterbunny17):

@AlexandervonHumboldt2 Please do not give out direct answers as doing so is against this site's Code of Conduct.

OpenStudy (mathstudent55):

\(y \div 2 \cdot 4 - 10 \div x + 6\) where \(x = 5\) and \(y = 16\) The first step is to replace x with 5 and y with 16. Copy the entire expression and make those substitutions. What do you get?

OpenStudy (alexandervonhumboldt2):

trhere is nothing to explain here order of operations you mean? no i listen to moderators because in COC if i did a mistake i need to listen moderarts forgt

OpenStudy (solomonzelman):

Lets take it step by step. x=5 y=16 \(\large\color{black}{y \div 2 \times 4 - 10 \div x + 6 }\) (you can use my code `\(\large\color{black}{y \div 2 \times 4 - 10 \div x + 6 }\) ` ) plug in 5 for x, and 16 for y first. Can you do that?